# Channel Twinning: An Enabler for Next-Generation Ubiquitous Wireless Connectivity

### v1

@@ III-C Lightweight Twin Modeling Module
Existing CT models typically train in a centralized manner. However, scalability and privacy challenges limits the CT deployment in ubiquitous connectivity. In this case, federated learning (FL) is viewed as an adequate and promising choice for lightweight CT because FL training is decentralized and does not require sharing data. Fig. 3 shows a lightweight CT training module based on FL. In the FL-enabled CT scheme, FL clients only need to upload local model parameters to collaboratively train a global CT model, thus reducing communication costs and privacy leakage risks.
The integration of FL and DT is not straightforward, as two practical issues need to be addressed as prerequisites. One is that participating in FL tasks can increase the energy costs of client devices. For this reason, self-interested sensor devices may be reluctant to participate in FL, which needs proper FL incentive mechanisms [15]. Clients participating in FL should meet their own energy constraints. For example, UAVs are energy-limited. In this sense, an incentive mechanism for FL-based CT should be designed to improve CT training efficiency while meeting the power constraints. To achieve such energy-efficient FL incentives, we need to quantify the contribution of clients to the global CT model convergence and design appropriate payment mechanisms. The other practical issue that needs to be addressed for the integration of FL and DT is the heterogeneity of local data on each client. This would require handling data distribution imbalances across clients. The reason is that personalized FL may be suitable for lightweight CT in the future.
@@ IV Case Studies: Intelligent Coverage and Lightweight Twin
Fig. 4: User association results produced based on the true channel, the twin channel, and the fitted PL model: (a) True channel gain; (b) twin channel gain; (c) channel gain predicted by fitted PL model; and (d) the achievable sum rate with different schemes. Each UE selects five APs for communications, and the associated APs are indicated by red dots.
Fig. 5: Federated twin of the channel gains: (a) The validation loss versus communication rounds, where the total number of training samples is 10,000, and these samples are randomly divided into three or five UEs for FL; and (b) MAE comparison of channel gain prediction across different models.
@@ IV-A CT-Based Intelligent Coverage
In ubiquitous connectivity networks, user association can dynamically adjust the connection of user equipment (UE) according to the channel status, in order to achieve continuous and efficient coverage. By user association, the optimal access points (APs) are selected for UE access, affecting network resource utilization. Dynamic adaptive user association can reduce latency or packet loss, avoid interruption, and save the transmit powers. However, the premise of dynamic adaptive user association is an accurate perception of full CSI.
@@
Cell-free communication, as a key ubiquitous connectivity technology, is taken as an example to illustrate the advantage of CT. In cell-free networks, obtaining the full CSI of involved nodes is challenged by limited pilots and high CSI exchange overhead. To address this, we propose a CT-based intelligent user association scheme in cell-free networks. Specifically, the proposed scheme is implemented via offline CT model training and online CSI query, as illustrated in Fig. 3(a). In the offline phase, dedicated measurement anchors are responsible for collecting channel measurements at various locations in the RoI and uploading measurements to the CT server for training. Once the CT model is obtained through offline training, it is stored in the twin repository. In the online phase, APs or UEs request the CT server to query the channel information of the node they attempt to communicate with.
To demonstrate the advantage of CT, we leverage the synthetic data generated by Wireless InSite [^2] https://www.remcom.com/wireless-insite-propagation-software software for model training [2]. Channel gain measurements, sampled at a uniform grid in the region shown in Fig. 4(a), serve as the ground truth.
@@
Based on the spatial interpolation concept, the trained CT model can capture the impact of environmental factors on channel gains. Fig. 4(b) shows a channel gain map given by a UE position, where the UE is indicated by a white square. Conventional user association is based only on the distance or path loss (PL) in practical cell-free systems. We derive the PL model with the same synthetic data by least squares fitting. As a benchmark, PL-based results are shown in Fig. 4(c). By comparing Figs. 4(a), 4(b), and 4(c), it can be observed that CT can efficiently perceive the environment, while the fitted PL model cannot reproduce the environment profile.
With the predicted channel gain, we select five APs from 20 distributed APs within the RoI for user association. Since the PL model does not grasp the environment semantics, it can only select APs according to the distance criterion. The channel gain comprises PL, shadowing, and multi-path effects. PL is determined by distance while shadowing and multi-path effect are affected by surroundings.
By contrast, CT can guide the UE in avoiding buildings, instead of directly selecting the nearest APs. Fig. 4(d) shows the sum rates achieved by different user association schemes, which confirms the benefits of CT for ubiquitous coverage. In addition, since open street maps [10] and user locations are reasonably easy to obtain, the CSI of any AP-UE link can be retrieved by CT. Thus, the dreadful CSI-sharing overhead can be avoided in ubiquitous connectivity.
@@ IV-B Federated Learning-Based Lightweight CT
In the FL-based CT training architecture, no dedicated measurement anchors are needed for data collection. Each mobile UE may be a potential measurement anchor. These incentive UEs can record their locations and associated channel measurements as they move into the RoI. Unlike centralized training, the UEs do not upload the data; rather, they first register their identities with the CT server’s client manager for subsequent FL training. When the CT server prepares to train a CT model, a group of registered UEs train their own CT models locally by using their recorded measurements. Then, they upload only the trained parameters to the server for model aggregation, thus obtaining the global CT model. The CT server sends the global model parameters to the registered clients to update the local models. After several rounds of interactions, the global model converges, and hence, the FL-based CT training is completed.
We select a portion of the channel gain dataset used in the previous subsection, randomly splitting and distributing it to FL clients. Fig. 5(a) shows the trained results. It is seen that more clients contribute to better validation loss performance, even if the total data samples remain the same. With more clients, the gradient noise of individual clients can be smoothed out. This indicates that increasing the number of clients can improve the federated twin performance, but raises communication overhead concerns.
Fig. 5(b) compares prediction performance among the centralized CT model, fitted PL model, and federated CT model in terms of mean absolute error (MAE). From the boxplots of MAE, the red solid line denotes the median MAE, the box represents the distribution range of the majority of MAE, and the black dots outside the box denote outliers. It is observed that the PL model exhibits high prediction errors overall, indicating poor prediction accuracy. Meanwhile, few outliers indicate prediction stability, as the channel gain produced by the PL model lacks shadowing and SSF compared to the true channel. Both CT models show better prediction accuracy. Even decentralized training-based CT model has superior accuracy to the PL model. Despite the fluctuations in prediction error of CT models, occasional outliers do not affect the overall performance of ubiquitous connectivity. Notably, the goal of CT is not to copy physical channel entities, but rather to satisfy network requirements by emulation.
@@ V-A1 CT-Empowered Non-Terrestrial Communications
Due to the high mobility of non-terrestrial systems, both satellite-to-ground and UAV-to-ground channels suffer from the non-negligible channel aging effect. Unlike the ground channel, UAV-to-ground channel gains do not show smooth changes with the movement of the UAVs. Due to the down-tilted antenna pattern, UAVs may easily fall into the antenna nulls of nearby BSs. Therefore, the UAVs must search for the best antenna sidelobe associated with the remote cell when switching cells instead of using the aforementioned distance criterion. In this case, CT can emulate the antenna gain and airframe shadowing [1] to achieve high-efficiency cell association in UAV networks.
@@ V-A2 CT-Empowered IoV Communications
One of the most immediate applications of CT is predictive beamforming in IoVs. CT-enabled beamforming must only send necessary control signaling without extra pilots [4]. On the other hand, RIS is regarded as an attractive low-power technology that can overcome dynamic blockage in IoVs. However, CSI acquisition and frequent beam alignment in high-mobility scenarios hinder the widespread application of RIS. Through the CT model dedicated to serving the RIS sites, adaptive beam width control and vehicle tracking can be further realized.
@@ V-A3 CT-Empowered Maritime Communications
CT is promising in achieving ubiquitous maritime connectivity. Note that the maritime communications environment has distinct features of sparse scatters compared to terrestrial networks. This motivates us to employ the environmental information to deploy CT models for network planning [3]. Channel interference can be acquired in advance to avoid the acquisition burden of full CSI. In the harsh underwater environment, underwater acoustic (UWA) channels suffer from spatio-temporal-frequency variability, limited bandwidth and severe noise and interference [1]. Channel estimation and data transmission becomes even more challenging in multi-node UWA communications. In this case, CT also brings unprecedented opportunities for UWA communications.
@@ V-B1 Data Deficiency
CT demands sufficient high-quality measurement data. However, channel measurement campaigns are time-consuming and expensive. In practice, only sparse measurements from limited points are available for CT training. For example, DL-based indoor propagation prediction [12] only utilizes channel measurements from 36 locations. Furthermore, some measurements may contain errors due to hardware defects. Therefore, for high-precision CT training, it is necessary to study the efficient utilization of synthetic and measurement data.
@@ V-B2 Personalized FL Mechanism of CT
Since FL clients only collect channel measurements from the local irregular areas, a simple federated averaging-based aggregation strategy can hardly be optimal. In particular, personalized FL with incentive mechanisms needs to consider the CT model contribution of each client to the entire environment map.
@@ V-B3 Malicious CT Deployment
Malicious nodes with CT ability may exist in networks. They can train the CT models based on local radio measurements or query the CT server for CSI at locations of interest. Therefore, once a malicious CT is able to recover CSI, wireless networks will face unparalleled security threats. Existing physical layer security and secret key schemes will be rendered ineffective.
@@ V-B4 High-Dimensional DTC Representation
While DTC is effective for wireless network optimization, high-dimensional channel matrices caused by large-scale MIMO or RIS raise computational and caching challenges for CT deployment. In this case, an equivalent channel compression representation scheme in CT is attractive for network optimization.
@@ VI Conclusion
From the perspective of the physical environment, a new channel acquisition paradigm, namely, CT technology, has been studied to support xG ubiquitous wireless communication networks. With multimodal sensing, CT can evolve to predict CSI that closely resembles actual channels. This article outlines recent CT-related activities with different conceptual names. Then, the concept, key features, and two case studies of CT have been discussed. To promote CT deployment in wireless networks, a modularized design architecture of CT has been presented. The architecture includes universal channel modeling of full scenarios, multimodal cooperative sensing, and FL-based lightweight CT training. New opportunities for CT in enabling technologies for ubiquitous connectivity and future research challenges have also been discussed.

### v3

@@ III-C Lightweight Twin Modeling Module
Existing CT models typically train in a centralized manner. However, scalability and privacy challenges limit the CT deployment in ubiquitous connectivity. In this case, federated learning (FL) is viewed as an adequate and promising choice for lightweight CT because FL training is decentralized and does not require sharing data. Fig. 2 shows a lightweight CT training module based on FL. In the FL-enabled CT scheme, FL clients only need to upload local model parameters to train a global CT model collaboratively, thus reducing communication costs and privacy leakage risks.
The integration of FL and DT is not straightforward, as two practical issues need to be addressed as prerequisites. First, participating in FL tasks can increase the energy costs of client devices. Self-interested devices may be reluctant to participate in FL. Thus, proper FL incentive mechanisms [14] are needed to encourage participation while considering energy constraints, especially for energy-limited devices like UAVs. To achieve such energy-efficient FL incentives, we need to quantify the contribution of clients to the global CT model convergence and design suitable payment mechanisms. Second, the heterogeneity of local data across clients requires handling data distribution imbalances, as personalized FL could become promising for lightweight CT in the future.
@@ IV Case Studies: Intelligent Coverage and Lightweight Twin
Fig. 3: User association results produced based on the ground-truth channel, the twin channel, and the fitted PL model.
@@ IV-A CT-Based Intelligent Coverage
In ubiquitous connectivity networks, user association can dynamically adjust the connection of user equipment (UE) according to channel conditions to ensure continuous and efficient coverage. By selecting optimal access points (APs) for UE access, user association can improve resource utilization, reduce packet loss, and prevent interruptions. However, accurate full CSI is essential for effective user association.
@@
Cell-free communication is taken as an example to illustrate the benefits of CT for ubiquitous connectivity. In cell-free networks, obtaining the full CSI of involved nodes is challenged by limited pilots and high CSI exchange overhead. To address this, we propose a CT-based intelligent user association scheme. Specifically, the scheme is implemented via offline CT model training and online CSI query, as illustrated in Fig. 2(a). During the offline phase, dedicated measurement anchors collect channel measurements in the RoI and upload measurements to the CT server for training. Once trained, the CT model is stored in the twin repository. During the online phase, UEs request the CT server to query the CSI of the node they attempt to communicate with.
We leverage the synthetic data generated by the WI software [8] to train a CT model. As shown in Fig. 3, the transceiver locations are selected from grid points with a uniform spacing of 8 meters. We then use a deep neural network with seven hidden layers to capture how the environment affects the channel gain distribution. The network takes four-dimensional transceiver coordinates as input and outputs the corresponding channel gains. For any given transmitter location, we perform SI [2] using the predicted values from 30 randomly selected receiver positions, enabling the generation of the channel gain maps.
@@
Fig. 3 shows channel gain maps with the UE marked by a white square. Conventional user association relies on the distance or path loss (PL) only. We fit a PL model with the same synthetic data by least squares fitting as a benchmark. By comparing the maps in Fig. 3, CT can efficiently perceive the environment, while the fitted PL model fails to reproduce the environment profile.
With the predicted channel gain, we select five out of 20 distributed APs within the RoI for user association. The PL model can only select APs according to the distance criterion, as it does not grasp the ESs. By contrast, CT helps the UE avoid obstacles rather than selecting the nearest APs. In addition, OpenStreetMap (OSM) data [9] can assist the CT model by generating synthetic data, particularly in scenarios where RF measurements might be sparse or unavailable, the CSI of any AP-UE link can be retrieved by CT.
Fig. 4: Validation loss of MSE versus communication rounds, where the total number of training samples is 10,000, and these samples are randomly divided between three or five UEs for FL.
Fig. 5: MAE comparison of channel gain prediction across different models.
@@ IV-B Federated Learning-Based Lightweight CT
In the FL-based CT training architecture, no dedicated measurement anchors are needed. Mobile UEs act as potential measurement anchors. Incentive UEs can record their locations and associated channel measurements as they move into the RoI. Unlike centralized training, the UEs first register their identities with the CT server’s client manager for subsequent FL training. When the CT server prepares to train a CT model, a group of registered UEs train their own CT models locally by using their recorded measurements. Then, they upload only the trained parameters to the server for model aggregation. The CT server updates the global model and sends model parameters back to the registered clients for local model updates. After several rounds of interactions, the global model converges, and hence, the FL-based CT training is completed.
A portion of the channel gain dataset was randomly split among FL clients. Fig. 4 shows the training loss of mean squared error (MSE). It is seen that more clients contribute to lower MSE loss, even with the same total data samples. This is because more clients help smooth out the gradient noise of individual clients and improve the federated twin performance, though communication overhead rises.
Fig. 5 compares the channel gain (dB) prediction capabilities in terms of mean absolute error (MAE). The yellow solid line denotes the median MAE, the box represents the distribution range of the majority of MAE, and the black dots outside the box denote outliers. The PL model exhibits high prediction errors overall, indicating poor prediction accuracy. Meanwhile, few outliers indicate prediction stability, as the channel gain produced by the PL model lacks shadowing compared to the real channel. Both CT models show better prediction accuracy. Despite the fluctuations in prediction error of CT models, occasional outliers do not affect the overall performance of ubiquitous connectivity.
In FL, the primary feedback overhead arises from the transmission of model updates from the UE to the BS. The feedback overhead of the 3GPP Type II codebook scales with the rank, and requires hundreds of bits (more than 500 bits for some system setups) [15]. Since the overhead of different setups varies greatly, it is difficult to pre-determine the required overhead. Thus, the BS allocates physical uplink shared channel (PUSCH) resources for CSI reporting according to the maximum overhead, which causes resource waste and reduced efficiency. In this case, the UE can utilize leftover PUSCH resources for FL model parameter uploads after completing CSI reporting, thus improving resource utilization. Additionally, once the FL model is trained for large time-scale invariant environments, it will significantly reduce the channel feedback overhead.
@@ V Convergence of CT and Ubiquitous Communications: Opportunities and Challenges
In addition to reducing pilot overhead and enhancing prediction accuracy, CT can benefit ubiquitous connectivity by offering enhanced dynamic system robustness [12], accelerated network development and testing [3], and improved resource management [4]. To unfold the potential of CT, this section presents some open issues concerning the convergence of CT and ubiquitous communications.
@@ V-A1 CT-Empowered Non-Terrestrial Communications
Due to the high mobility of non-terrestrial systems, both satellite-to-ground and UAV-to-ground channels suffer from the non-negligible channel aging effect. Unlike the ground channel, UAV-to-ground channel gains do not show smooth changes with the movement of the UAVs. Due to the down-tilted antenna pattern, UAVs may easily fall into the antenna nulls of nearby BSs. Therefore, the UAVs cannot use the distance criterion to choose the cell. In this case, CT can emulate the antenna gain and airframe shadowing [1, 4] to achieve high-efficiency cell association in UAV networks.
@@ V-A2 CT-Empowered IoV Communications
One of the most immediate applications of CT is predictive beamforming in IoV. CT-enabled beamforming requires only necessary control signaling without extra pilots [5]. On the other hand, RIS is regarded as a low-power solution to dynamic blockage issues in IoV. However, CSI acquisition and frequent beam alignment in high-mobility scenarios hinder the widespread application of RIS. Through the CT model dedicated to serving the RIS sites, adaptive beam width control and vehicle tracking can be further realized.
@@ V-A3 CT-Empowered Maritime Communications
Since the maritime communications environment has distinct features of sparse scatters compared to terrestrial networks, CT is expected to enhance maritime connectivity by employing environmental information for network planning [3]. Channel interference can be acquired in advance to avoid the acquisition burden of full CSI. In the harsh underwater environment, underwater acoustic (UWA) channels suffer from spatio-temporal-frequency variability, limited bandwidth and severe noise and interference [1]. Channel estimation and data transmission become even more challenging in multi-node UWA communications. In this case, CT also brings unprecedented opportunities for UWA communications.
@@ V-B1 Data Deficiency
CT demands sufficient high-quality measurement data, but channel measurement campaigns are costly and time-consuming. In practice, only sparse measurements from limited points are available. For example, DL-based indoor propagation prediction [11] only uses measurements from 36 locations. Furthermore, some measurements may contain errors due to hardware defects. To achieve high-precision CT, it is necessary to study the efficient utilization of synthetic and measurement data.
@@ V-B2 Personalized FL Mechanism of CT
Since FL clients only collect channel measurements from the local irregular areas, a simple federated averaging-based aggregation strategy can hardly be optimal. In particular, personalized FL with incentive mechanisms [14] needs to consider the CT model contribution of each client to the entire environment map.
@@ V-B3 Malicious CT Deployment
Malicious nodes with CT ability may exist in networks. They may access CSI at any location of interest. Therefore, once a malicious CT can recover CSI, wireless networks will face unparalleled security threats. Existing physical layer security and secret key schemes will be rendered ineffective.
@@ V-B4 High-Dimensional DTC Representation
While DTC is effective for wireless network optimization, high-dimensional channel matrices caused by large-scale MIMO or RIS raise computational and caching challenges for CT deployment. In this case, an equivalent channel compression representation [7] scheme in CT is attractive for network optimization.
@@ VI Conclusion
From the perspective of the physical environment, a new CA paradigm, CT technology, has been studied to support xG ubiquitous wireless communication networks. With multimodal sensing, CT can evolve to predict CSI that closely resembles actual channels. This article outlines recent CT-related activities with different conceptual names. Then, the concept, key features, and two case studies of CT have been discussed. To promote CT deployment in wireless networks, a modularized design architecture of CT has been presented. The architecture includes versatile model configuration of full scenarios, multimodal cooperative sensing, and FL-based lightweight CT training. New opportunities for CT in enabling technologies for ubiquitous connectivity and future research challenges have also been discussed.